# cryoSENSE: Compressive Sensing Enables High-throughput Microscopy with Sparse and Generative Priors on the Protein Cryo-EM Image Manifold

### v2

@@ 1 Introduction
Recovering images from compressed measurements is a fundamentally ill-posed inverse problem, as infinitely many candidate images can be consistent with the same compressed observations. To address this challenge, cryoSENSE employs two complementary strategies [10, 11, 12]. First, cryoSENSE imposes handcrafted priors that constrain the solution space [13], particularly sparsity in predefined transform domains (e.g., wavelets or discrete cosine transform bases) [14], or regularization schemes such as total variation, which promote piecewise smoothness [15]. Such priors have demonstrated effectiveness across diverse imaging modalities, including magnetic resonance imaging [16], infrared and visible-wavelength imaging [17], four-dimensional scanning transmission electron microscopy [18, 19], and fluorescence microscopy [20, 21]. Second, cryoSENSE learns data-driven priors directly from experimental images using generative models [22, 23, 24, 25, 26]. In particular, it leverages denoising diffusion probabilistic models (DDPMs) [27, 28], replacing the assumption of sparsity in a fixed basis with the assumption that images lie on a low-dimensional manifold [29, 30, 31, 32, 33]. By learning to reverse a stochastic noising process trained on real cryo-EM data, diffusion models are guided to generate high-quality reconstructions from incomplete or corrupted measurements [22, 23, 24, 34, 35].
@@ 2 Method
In cryoSENSE, we solve the inverse problem of recovering a cryo-EM image $\mathbf{x}^{*}\in\mathbb{R}^{n}$ from noisy, compressed measurements $\mathbf{y}=\mathcal{A}(\mathbf{x}^{*})+\bm{\eta}\in\mathbb{R}^{m}$, where $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a known linear projection operator (in pixel or Fourier space) representing the acquisition of $m\ll n$ masked images and $\bm{\eta}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ represents additive white Gaussian noise (Fig. 2a). To solve this ill-posed problem, we assume two different priors over cryo-EM images: i) sparsity in a pre-defined basis and ii) living on a low-dimensional manifold $\mathcal{M}$ learned from the EMPIAR data. In the former case, we use proximal gradient descent, and in the latter, we do posterior sampling on a denoising diffusion model to recover the protein image $\mathbf{x}^{*}$.
@@ 2.3 Image Recovery with Generative Prior
$d\mathbf{x}={}$ $\biggl[-\frac{\beta_{t}}{2}\,\mathbf{x}-\beta_{t}\Bigl(\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})+\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})\Bigr)\biggr]dt+\sqrt{\beta_{t}}\,d\bar{\mathbf{w}}.$ (7)
@@
where $\sigma_{t}^{2}$ is a hyperparameter. The denoised estimate $\widehat{\mathbf{x}}_{0}$ allows us to compute $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\widehat{\mathbf{x}}_{0})$. Assuming $\mathcal{A}(\mathbf{x}_{0})$ is differentiable, this enables us to approximate $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})$ as $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})\simeq\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\widehat{\mathbf{x}}_{0})$ [24]. Assuming $\bm{\eta}$ is Gaussian noise with variance $\sigma^{2}$, $p(\mathbf{y}|\mathbf{x}_{0})$ takes the form of:
$p(\mathbf{y}|\mathbf{x}_{0})$ $\sim\mathcal{N}\!\left(\mathbf{y}|\mathcal{A}(\mathbf{x}_{0}),\,\sigma^{2}\mathbf{I}\right)=\frac{1}{\sqrt{(2\pi)^{n}\sigma^{2n}}}\exp\!\left(-\frac{\|\mathcal{A}(\mathbf{x}_{0})-\mathbf{y}\|_{2}^{2}}{2\sigma^{2}}\right).$ (9)
@@
$\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})$ $\simeq\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\widehat{\mathbf{x}}_{0})=-\frac{1}{\sigma^{2}}\,\nabla_{\mathbf{x}_{t}}\bigl\|\mathcal{A}(\widehat{\mathbf{x}}_{0})-\mathbf{y}\bigr\|_{2}^{2}.$ (10)
This approximation allows cryoSENSE to replace the term $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})$ with $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\widehat{\mathbf{x}}_{0})$ and facilitate efficient posterior sampling for cryo-EM image reconstruction of $\mathbf{x}^{*}$. To realize posterior sampling efficiently over a fixed $T$ number of steps, cryoSENSE integrates Nesterov-accelerated gradients to steer the reverse diffusion process [58, 59, 60]. For more details on the implementation of cryoSENSE, we refer to Appendices 5 and 6.
3 Experimental Results
Prior $K$ 22 44 88 $16$ $32$
$C$ 4 2 1.3 1 16 2.7 1.5 1 10.7 2.6 1.5 1 10.2 2.5 1.4 1 10 2.5 1.4 1
3.1 Experimental Setup
Datasets. We evaluate cryoSENSE on five publicly available cryo-EM datasets: EMPIAR-10076 [61] (ribosomal complex), EMPIAR-10648 [62] (protein-ligand complex), EMPIAR-10166 [63] (protein complex responsible for cellular homeostasis), EMPIAR-11526 [64] (ribosomal subunit in E. coli), and EMPIAR-10786 [65] (substance P-neurokinin receptor G protein complex). These datasets comprise single-particle cryo-EM images with corresponding Contrast Transfer Function (CTF) and pose metadata. All EMPIAR datasets are collected at $\times128$ resolution, with the exception of EMPIAR-10648 experiments, which are collected at $\times256$ resolution. Full dataset details and preprocessing protocols are provided in Appendix 7. In the main text, we include results from EMPIAR-10076 and EMPIAR-10648, and leave the remainder of the proteins in the Appendix.
Evaluation metrics. To assess image reconstruction quality, we use the Structural Similarity Index Measure (SSIM) [66] and Learned Perceptual Image Patch Similarity (LPIPS) [67], which provide complementary measures of structural and perceptual similarity between reconstructed and ground-truth images. For 3D volume evaluation, we backproject images using the provided particle poses to reconstruct three volumes: (1) the original images, (2) cryoSENSE-sparse prior with DCT, which achieved the best overall performance among all sparse priors across our image reconstruction experiments (see Appendix 10), and (3) cryoSENSE-generative prior reconstructed images. We then compute the achieved resolution (in Å) using Fourier Shell Correlation (FSC) at the 0.143 cutoff [68] with no mask, and assess real-space similarity using the volume-to-volume correlation coefficient (VC) via the measure correlation command in UCSF ChimeraX [69].
Measurement process. In our experiments, we implement the measurement operator $\mathcal{A}$ through both pixel-space masking and Fourier-space masking. We define the compression factor as $C=n/m\geq 1$, where a higher $C$ corresponds to fewer acquired measurements $m$ taken. For pixel-space masking, we linearly project $\mathbf{x}$ through a set of $b$ binarized random masks. For each masked image, we apply a non-overlapping kernel-wise summation over patches of size $K\times K$ across the entire image. This operation reduces spatial resolution and produces low-resolution measurements $\mathbf{y}$ of dimension $m=bn/K^{2}$, resulting in a compression factor of $C=K^{2}/b$. For Fourier-space masking, $\mathcal{A}$ first takes the Fourier transform of $\mathbf{x}$ and then applies a single binary mask to subsample $m$ total Fourier coefficients, such that $m=n/C$. To choose which Fourier coefficients to subsample, we explore three distinct masking strategies: (1) uniform subsampling, (2) annular ring sampling with a low-frequency bias, and (3) radial spoke sampling. For all Fourier-space masking 3D volume reconstructions, we use uniform subsampling, which is the best-performing masking strategy among all experiments. Full details about the experiments can be found in Appendices 8 to 10.
@@ 3.2 High-fidelity Reconstruction and Fundamental Acquisition Limits
We first benchmark cryoSENSE’s ability to reconstruct cryo-EM images from highly compressed measurements on EMPIAR-10076. We evaluate both pixel-space masking, across five downsampling levels $K\in\{2,4,8,16,32\}$ and varying $C$, and Fourier-space masking, varying $C$. The compressed measurements used for all experiments were obtained without noise. In Appendix 10, we additionally repeat the reconstruction process under additive Gaussian noise and show that moderate ($\sigma^{2}=0.1$) noise minimally impacts reconstruction quality.
@@
Table 3: Reconstruction performance on EMPIAR-10076 with pixel-space and Fourier-spacing masking, benchmarking FSC resolution values (Å) and real-space volume correlations with the ground-truth structure (VC). Compression factors $C$ are rounded to the nearest tenth. Ground truth structure has a resolution of 6.65 Å.
@@
Prior $K$ $C$ Å $\downarrow$ VC $\uparrow$ $C$ Å $\downarrow$ VC $\uparrow$
Sparse - DCT 2 4 9.53 0.86 10 10.75 0.87
4 2 7.35 0.91 5 8.38 0.95
8 1.5 6.65 0.95 3.3 7.35 0.97
16 1.3 7.35 0.82 2.5 6.65 0.98
32 1.4 14.46 0.76 2 6.65 0.99
Gen - DDPM 2 4 9.98 0.83 10 11.03 0.77
4 2 8.38 0.80 5 9.11 0.91
8 1.5 8.06 0.81 3.3 7.62 0.96
16 1.3 6.99 0.97 2.5 6.76 0.98
32 1.4 7.91 0.93 2 6.65 0.99
We additionally assess downstream structural fidelity by reconstructing 3D volumes with sparse priors (DCT) and generative priors (DDPM) using known particle poses (Table 3). In pixel-space masking, following our 2D analysis, we select $C$ for each $K$ based on achieving an LPIPS threshold of approximately 0.15 with DDPM. DCT notably recovers ground-truth resolution at $C=1.45$, while DDPM provides better reconstructions at large values of $K$. Similar to the 2D image case, in Fourier-space masking, both methods maintain high fidelity at compression factors up to $C=2.5$, with DCT outperforming DDPM across all compression factors. These results highlight the flexibility and robustness of cryoSENSE in masking strategies and compression and provide a powerful strategy for addressing cryo-EM throughput.
@@ 3.4 Atomic Model Recovery via ModelAngelo
To assess the structural accuracy of atomic models obtained from cryoSENSE reconstructions, we analyze atomic structures generated from cryo-EM 3D volumes of EMPIAR-10648 using ModelAngelo [71] (see Appendix 9.3 for inference details). We evaluate three cases: (1) cryo-EM 3D volume reconstructed from the validation particle set (Original) and cryoSENSE pixel-space masking reconstructions after $K=2$ downsampling with $C=1.33$ using (2) DCT and (3) DDPM priors. For all cases, atomic model building is performed using the corresponding cryo-EM 3D volumes and the FASTA sequence of Protein Data Bank (PDB) entry 6ttf, representing the sequence associated with EMPIAR-10648. Atomic models for each condition are colored by the ModelAngelo per-residue prediction confidence scores (Fig. 4a). The mean per-residue confidence scores are similarly high for all three conditions, with Original, DCT, and DDPM achieving 67.3 $\pm$ 33.5, 64.71 $\pm$ 33.82, 62.5 $\pm$ 33.4, respectively. We next examine the resolutions of the underlying 3D maps using FSC (Fig. 4b). The three volumes achieve FSC (tight mask) resolutions of 3.73 Å (Original), 3.78 Å (DCT), and 3.82 Å (DDPM), indicating that cryoSENSE retains sufficient structural detail to support reliable atomic model building, with the DCT prior providing the marginally higher-resolution map.
@@
Together, these results show that cryoSENSE supports accurate atomic model recovery from compressed measurements. At $K=2$ downsampling with $C=1.33$, the DCT prior performs slightly better overall, achieving a higher map resolution (3.78 Å vs. 3.82 Å), higher mean confidence scores (64.71 vs. 62.5), and more matched chains (58 vs. 42) than the DDPM prior.

### v3

@@ 1 Introduction
Recovering images from compressed measurements is a fundamentally ill-posed inverse problem, as infinitely many candidate images can be consistent with the same compressed observations. To address this challenge, cryoSENSE employs two complementary strategies [25, 14, 13]. First, cryoSENSE imposes handcrafted priors that constrain the solution space [84], particularly sparsity in predefined transform domains (e.g., wavelets or discrete cosine transform bases) [62], or regularization schemes such as total variation, which promote piecewise smoothness [79]. Such priors have demonstrated effectiveness across diverse imaging modalities, including magnetic resonance imaging [50], infrared and visible-wavelength imaging [5], four-dimensional scanning transmission electron microscopy [57, 56, 65], and fluorescence microscopy [51, 12]. Second, cryoSENSE learns data-driven priors directly from experimental images using generative models [15, 81, 17, 64, 74]. In particular, it leverages denoising diffusion probabilistic models (DDPMs) [35, 24], replacing the assumption of sparsity in a fixed basis with the assumption that images lie on a low-dimensional manifold [10, 39, 6, 28, 46]. By learning to reverse a stochastic noising process trained on real cryo-EM data, diffusion models are guided to generate high-quality reconstructions from incomplete or corrupted measurements [15, 81, 17, 73, 49].
@@ 2 Method
In cryoSENSE, we solve the inverse problem of recovering a cryo-EM image $\mathbf{x}^{*}\in\mathbb{R}^{n}$ from noisy, compressed measurements $\mathbf{y}=\mathcal{A}(\mathbf{x}^{*})+\boldsymbol{\eta}\in\mathbb{R}^{m}$, where $\mathcal{A}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ is a known linear projection operator (in pixel or Fourier space) representing the acquisition of $m\ll n$ masked images and $\boldsymbol{\eta}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ represents additive white Gaussian noise (Fig. 2a). To solve this ill-posed problem, we assume two different priors over cryo-EM images: i) sparsity in a pre-defined basis and ii) living on a low-dimensional manifold $\mathcal{M}$ learned from the EMPIAR data. In the former case, we use proximal gradient descent, and in the latter, we do posterior sampling on a denoising diffusion model to recover the protein image $\mathbf{x}^{*}$.
@@ 2.3 Image Recovery with Generative Prior
$d\mathbf{x}={}$ $\biggl[-\frac{\beta_{t}}{2}\,\mathbf{x}-\beta_{t}\Bigl(\nabla_{\mathbf{x}_{t}}\log p(\mathbf{x}_{t})$ (7)
$\qquad\quad{}+\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})\Bigr)\biggr]dt+\sqrt{\beta_{t}}\,d\bar{\mathbf{w}}.$
@@
where $\sigma_{t}^{2}$ is a hyperparameter. The denoised estimate $\widehat{\mathbf{x}}_{0}$ allows us to compute $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\widehat{\mathbf{x}}_{0})$. Assuming $\mathcal{A}(\mathbf{x}_{0})$ is differentiable, this enables us to approximate $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})$ as $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})\simeq\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\widehat{\mathbf{x}}_{0})$ [17]. Assuming $\boldsymbol{\eta}$ is Gaussian noise with variance $\sigma^{2}$, $p(\mathbf{y}|\mathbf{x}_{0})$ takes the form of:
$p(\mathbf{y}|\mathbf{x}_{0})$ $\sim\mathcal{N}\!\left(\mathbf{y}|\mathcal{A}(\mathbf{x}_{0}),\,\sigma^{2}\mathbf{I}\right)$ (9)
$=\frac{1}{\sqrt{(2\pi)^{n}\sigma^{2n}}}\exp\!\left(-\frac{\|\mathcal{A}(\mathbf{x}_{0})-\mathbf{y}\|_{2}^{2}}{2\sigma^{2}}\right).$
@@
$\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})$ $\simeq\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\widehat{\mathbf{x}}_{0})$ (10)
$=-\frac{1}{\sigma^{2}}\,\nabla_{\mathbf{x}_{t}}\bigl\|\mathcal{A}(\widehat{\mathbf{x}}_{0})-\mathbf{y}\bigr\|_{2}^{2}.$
This approximation allows cryoSENSE to replace the term $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\mathbf{x}_{t})$ with $\nabla_{\mathbf{x}_{t}}\log p(\mathbf{y}|\widehat{\mathbf{x}}_{0})$ and facilitate efficient posterior sampling for cryo-EM image reconstruction of $\mathbf{x}^{*}$. To realize posterior sampling efficiently over a fixed $T$ number of steps, cryoSENSE integrates Nesterov-accelerated gradients to steer the reverse diffusion process [80, 47, 42]. For more details on the implementation of cryoSENSE, we refer to the supplemental material.
@@
3 Experimental Results
3.1 Experimental Setup
Datasets. We evaluate cryoSENSE on five publicly available cryo-EM datasets: EMPIAR-10076 [22] (ribosomal complex), EMPIAR-10648 [70] (protein-ligand complex), EMPIAR-10166 [34] (protein complex responsible for cellular homeostasis), EMPIAR-11526 [76] (ribosomal subunit in E. coli), and EMPIAR-10786 [33] (substance P-neurokinin receptor G protein complex). These datasets comprise single-particle cryo-EM images with corresponding Contrast Transfer Function (CTF) and pose metadata. All EMPIAR datasets were collected at $\times128$ resolution, with the exception of EMPIAR-10648 experiments, which was collected at $\times256$ resolution. Full dataset details and preprocessing protocols are provided in the supplementary material. In the main text, we include results from EMPIAR-10076 and EMPIAR-10648, and leave the remainder of the proteins in the supplemental material.
Evaluation metrics. To assess image reconstruction quality, we use the Structural Similarity Index Measure (SSIM) [82] and Learned Perceptual Image Patch Similarity (LPIPS) [86], which provide complementary measures of structural and perceptual similarity between reconstructed and ground-truth images. Peak-signal-to-noise-ratio (PSNR) results are provided in the supplementary material. For 3D volume evaluation, we backproject images using the provided particle poses to reconstruct three volumes: (1) the original images, (2) cryoSENSE-sparse prior with DCT, which achieved the best overall performance among all sparse priors across our image reconstruction experiments (see supplementary material), and (3) cryoSENSE-generative prior reconstructed images. We then compute the achieved resolution (in Å) using Fourier Shell Correlation (FSC) at the 0.143 cutoff [66] with no mask, and assess real-space similarity using the volume-to-volume correlation coefficient (VC) via the measure correlation command in UCSF ChimeraX [53].
Measurement process. In our experiments, we implement the measurement operator $\mathcal{A}$ through both pixel-space masking and Fourier-space masking. We define the compression factor as $C=n/m\geq 1$, where a higher $C$ corresponds to fewer acquired measurements $m$ taken. For pixel-space masking, we linearly project $\mathbf{x}$ through a set of $b$ binarized random masks. For each masked image, we apply a non-overlapping kernel-wise summation over patches of size $K\times K$ across the entire image. This operation reduces spatial resolution and produces low-resolution measurements $\mathbf{y}$ of dimension $m=bn/K^{2}$, resulting in a compression factor of $C=K^{2}/b$. For Fourier-space masking, $\mathcal{A}$ first takes the Fourier transform of $\mathbf{x}$ and then applies a single binary mask to subsample $m$ total Fourier coefficients, such that $m=n/C$. To choose which Fourier coefficients to subsample, we explore three distinct masking strategies: (1) uniform subsampling, (2) annular ring sampling with a low-frequency bias, and (3) radial spoke sampling. For all Fourier-space masking 3D volume reconstructions, we use uniform subsampling, which is the best-performing masking strategy among all experiments. Full details about the experiments can be found in the supplemental material.
We first benchmark cryoSENSE’s ability to reconstruct cryo-EM images from highly compressed measurements on EMPIAR-10076. We evaluate both pixel-space masking, across five downsampling levels $K\in\{2,4,8,16,32\}$ and varying $C$, and Fourier-space masking, varying $C$. The cryo-EM images from EMPIAR-10076 already contain experimental noise. For the main experiments, the compressed measurements are generated from these images without adding any additional measurement noise after applying the measurement operator. In the supplementary material, we additionally repeat the reconstruction process under additive Gaussian measurement noise and show that moderate ($\sigma^{2}=0.01$) noise minimally impacts reconstruction quality. We further provide a detailed discussion of the measurement noise model and evaluate noise robustness across different measurement SNR levels.
@@ 3.2 High-fidelity Reconstruction and Fundamental Acquisition Limits
Table 3: Reconstruction performance on EMPIAR-10076 with pixel-space and Fourier-space masking, benchmarking FSC resolution values (Å), real-space volume correlations with the ground-truth structure (VC), and pose estimation accuracy, reporting median angular error (AE, in degrees) and median translational shift error (TSE, in Å). Compression factors $C$ are rounded to the nearest tenth. Ground truth structure has a resolution of 6.65 Å.
@@
Prior $K$ $C$ Å $\downarrow$ VC $\uparrow$ AE TSE $C$ Å $\downarrow$ VC $\uparrow$ AE TSE
Sparse - DCT 2 4 9.53 0.86 2.6 1.4 10 10.75 0.87 130.9 28.6
4 2 7.35 0.91 1.4 0.8 5 8.38 0.95 124.2 22.0
8 1.5 6.65 0.95 1.4 0.7 3.3 7.35 0.97 92.2 12.7
16 1.3 7.35 0.82 2.8 1.4 2.5 6.65 0.98 4.9 2.3
32 1.4 14.46 0.76 129.8 24.3 2 6.65 0.99 1.9 0.9
Gen - DDPM 2 4 9.98 0.83 3.1 1.6 10 11.03 0.77 128.8 22.1
4 2 8.38 0.80 1.9 1.0 5 9.11 0.91 116.4 16.6
8 1.5 8.06 0.81 4.3 2.2 3.3 7.62 0.96 72.7 8.8
16 1.3 6.99 0.97 1.7 0.9 2.5 6.76 0.98 3.5 2.5
32 1.4 7.91 0.93 4.6 2.3 2 6.65 0.99 1.8 0.9
We additionally assess downstream structural fidelity by reconstructing 3D volumes with sparse priors (DCT) and generative priors (DDPM) using known particle poses (Table 3). In pixel-space masking, following our 2D analysis, we select $C$ for each $K$ based on achieving an LPIPS threshold of approximately 0.15 with DDPM. DCT notably recovers ground-truth resolution at $C=1.45$, while DDPM provides better reconstructions at large values of $K$. Similar to the 2D image case, in Fourier-space masking, both methods maintain high fidelity at compression factors up to $C=2.5$, with DCT outperforming DDPM across all compression factors. We further evaluate pose estimation robustness by performing homogeneous refinement in CryoSPARC [61] on cryoSENSE reconstructions (Table 3). In pixel-space masking, poses remain accurate (median angular error $5^{\circ}$, shift error <2.5<2.5 Å) up to $K=16$, with DDPM maintaining accurate poses at $K=32$ whereas DCT does not. In Fourier-space masking, poses remain accurate up to $C=2.5$, consistent with the compression limits identified above. These results highlight the flexibility and robustness of cryoSENSE in masking strategies and compression and provide a powerful strategy for addressing cryo-EM throughput.
@@ 3.4 Atomic Model Recovery via ModelAngelo
To assess the structural accuracy of atomic models obtained from cryoSENSE reconstructions, we analyze atomic structures generated from cryo-EM 3D volumes of EMPIAR-10648 using ModelAngelo [40] (see supplementary material for inference details). We evaluate three cases: (1) cryo-EM 3D volume reconstructed from the validation particle set (Original) and cryoSENSE pixel-space masking reconstructions after $K=2$ downsampling with $C=1.33$ using (2) DCT and (3) DDPM priors. For all cases, atomic model building is performed using the corresponding cryo-EM 3D volumes and the FASTA sequence of Protein Data Bank (PDB) entry 6ttf, representing the sequence associated with EMPIAR-10648. Atomic models for each condition are colored by the ModelAngelo per-residue prediction confidence scores (Fig. 4a). The mean per-residue confidence scores are similarly high for all three conditions, with Original, DCT, and DDPM achieving 67.3 $\pm$ 33.5, 64.71 $\pm$ 33.82, 62.5 $\pm$ 33.4, respectively. We next examine the resolutions of the underlying 3D maps using FSC (Fig. 4b). The three volumes achieve FSC (tight mask) resolutions of 3.73 Å (Original), 3.78 Å (DCT), and 3.82 Å (DDPM), indicating that cryoSENSE retains sufficient structural detail to support reliable atomic model building, with the DCT prior providing the marginally higher-resolution map.
@@
Together, these results show that cryoSENSE supports accurate atomic model recovery from compressed measurements. At $K=2$ downsampling with $C=1.33$, the DCT prior performs slightly better overall, achieving a higher map resolution (3.78 Å vs. 3.82 Å), higher mean confidence scores (64.71 vs. 62.5), and more matched chains (58 vs. 42) than the DDPM prior. In the supplementary material, we additionally compare cryoSENSE against DMPlug [81], a diffusion-based super-resolution baseline that does not enforce measurement consistency with compressed measurements, and show that cryoSENSE outperforms it across all downstream tasks.